OpenStudy (anonymous):

for 3/x-1=10, is it x=1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if it is \(\frac{3}{x-1}=10\) then if \(x=1\) you would have \(\frac{3}{1-1}=\frac{3}{0}\) which is undefined, and certainly not equal to ten

OpenStudy (anonymous):

yes it i the equation i am working on.. solve graphically, so i multiplied both sides by x-1, then divide by ....10?

OpenStudy (anonymous):

I would cross multiply.

OpenStudy (anonymous):

no idea... i feel like im missing a number

OpenStudy (anonymous):

Are you expecting an integer answer?

OpenStudy (anonymous):

\[\frac{3}{x-1} = \frac{10}{1}\]

OpenStudy (anonymous):

says to simplify answer and use commas as needed-??- or other answer is "all real numbers"

OpenStudy (anonymous):

Start by cross multiplying. What do you get?

OpenStudy (anonymous):

30/x-1

OpenStudy (anonymous):

Do you know how to cross multiply? 3 * 1 = 10(x-1)

OpenStudy (anonymous):

oh thats right..

OpenStudy (anonymous):

guess not...:(

OpenStudy (anonymous):

OK, let me draw a picture (a picture is worth 1000 words, right?)

OpenStudy (anonymous):



OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

\[\frac{x}{y} = \frac{a}{b}\] \[x*b = y*a\]

OpenStudy (anonymous):

So, what did you get when you cross multiplied?

OpenStudy (anonymous):

3/10x-10

OpenStudy (anonymous):

You seem to be "hooked on fractions". See if this helps... \(\frac{3}{x-1} = \frac{10}{1}\) 3 * 1 = 10(x-1)